Shift Summary for Owl 1/24/00


Subject: Shift Summary for Owl 1/24/00
From: Vivian O'Dell (odell@fnal.gov)
Date: Fri Feb 04 2000 - 15:39:17 CST


              Shift Summary for owl shift, 24/Jan/99

On Shift: A. Bellavance (Walk-thru)

Hours of beam delivered at all intensities: 0 as expected
========================================================================
Description of Shift:

time comment
------------------------------------------------------------------------
02:00 Arrive at a quiet hall. Do gas checklist. Everything seems
          fine. There is an unusual light blinking on the south alarm
          panel: "Supply Fan SF-102 off". 20 minutes later it is no
          longer blinking. Must have corrected itself.

05:15 Another peaceful walk-thru (included downstairs this time). I
          notice the processes on atev1 have not updated since 11am
          yesterday. I kill both the chamber HV current window and the
          large DC/TRD window hoping they will re-start themselves. The
          DC current plot restarts; the other one doesn't. Do an SDAQ
          fresh_start. The large DC/TRD window comes back but isn't
          updating. Ah, it's complaining about no spill signal. Guess I
          should have just left it alone.

========================================================================
Items to Follow up on Next Shift

-The DC high voltages look like they are still on. Should they be off?

_______
Angela Bellavance work: (630) 840-2801
bellavan@rice.edu home: (630) 548-3412
http://www.owlnet.rice.edu/~bellavan/ page: (630) 722-6861
"It has yet to be proven that intelligence has any survival value."
                                - Arthur C. Clarke



This archive was generated by hypermail 2b27 : Fri Feb 04 2000 - 15:39:20 CST